# A Survey on Integrated Sensing, Communication, and Computation

### v1

@@ I-C Organization
  • Section V introduces three types of signal designs for ISCC, i.e., single-/dual-/triple-functional signal designs. Particularly, triple-functional signal design refers to using one signal to perform three functions of sensing, communication, and computation.
  • In Section VI, the network resource management paradigms for ISCC are discussed. One is joint RRM for task coexistence, which refers to resource allocation among different tasks of sensing, communication, and computation to achieve multiple goals simultaneously. The other is task-oriented ISCC resource allocation, which refers to allocating network resources among the sensing, communication, and computation modules of a task to achieve its customized goal.
  • The implementation of ISCC in future advanced networks is investigated in Section VII, including ISCC for digital twin enabled wireless networks, computing power networks supported ISCC, SAGINs supported ISCC.
  • Section VIII concludes the paper.
@@ II-A Overview
Emerging intelligent services over wireless networks often encompass both communication and computation processes. In conventional wireless networks, the communication and computation processes are separately designed. For instance, service tasks are typically executed at a central server with high computing and storage capabilities, where the central server first collects the data from geographically dispersed devices for task-related computation and then feeds back the computation results to corresponding devices for fulfilling the service requirements. With the proliferation of wirelessly connected devices, the communication burden of bandwidth-limited wireless networks significantly increases and the computation burden of the central server is also augmented due to massive volumes of data to be processed, making it inefficient to handle diverse service requirements in a centralized manner. Such a separate design ignores the inherent connection between the transmitted data and subsequent service tasks, leading to inefficient resource utilization and in turn degrading the system performance.
By treating the communication and computation processes as a unified module, ICC emerges as a promising technique to enhance the system performance, where the allocation of communication and computation resources are jointly optimized to achieve a common objective, e.g., energy/delay minimization and computation accuracy maximization, thereby making full use of limited resources to improve the system performance. In this section, we provide an overview of the existing studies on ICC, focusing on the efficient design of joint communication and computation, which is illustrated by presenting two promising examples, e.g., MEC and AirComp, as shown in Fig. 2 and listed below.
@@ II-B Mobile Edge Computing
In the past few years, computation tasks over wireless networks have been typically accomplished by resorting to cloud computing, which requires the large amount of data generated by devices to be collected and transmitted to a centralized server with high storage and computing capabilities, so as to perform data analysis and computation for meeting various requirements of different devices. However, with the rapid development of Internet of Things (IoT), the number of connected devices dramatically increases, leading to a severe communication burden of the wireless network and high computing pressure of the central server. Therefore, due to limited radio resources, cloud computing is no longer suitable for latency-sensitive services with massive connectivities in future wireless networks. To address the above issues, MEC is emerging as a new computing paradigm thanks to the advancement of microchip computing power, which enables the edge server and mobile devices to execute computing tasks with their own computing and storage capabilities at the network edge [12].
@@
A typical MEC system consists of edge servers and mobile devices, where edge servers are generally co-located with access points (APs) and deployed in proximity to mobile devices. Herein, the mobile devices are allowed to offload the computational tasks to edge servers for executing the tasks with high computing pressure and low latency requirements. To unleash the potential of MEC, numerous studies have explored task offloading design in MEC systems with single/multiple edge server(s) and multiple users. Specifically, the system design of MEC lies in the following aspects.
@@
  • Joint Resource Allocation: Even though the edge servers are capable of providing strong computing power to fulfill various service requirements for different devices, they still confront high computation and communication burdens when there are massive tasks from multiple devices that need to be accomplished due to the limited radio and computing resources of the MEC system. To overcome the above challenge, joint communication and computation resource allocation is required to enhance the system performance in terms of different objectives, e.g., execution delay and energy consumption. For instance, the authors in [50] propose a task offloading framework for MEC in software-defined ultra-dense networks, where the task placement and resource allocation are jointly optimized to minimize the average task duration under the given battery energy capacity of each device. Besides, the authors in [92] develop a SAGIN-assisted MEC for computing offloading in remote areas lacking conventional edge/cloud infrastructure, where a learning-based joint resource allocation and task scheduling approach is proposed to reduce the execution delay. To mitigate the overhead for collecting the global information to enable centralized system design, the authors in [93] propose an efficient decentralized algorithm for the task offloading in a dense wireless network with selfish mobile devices. By using game-theoretical analysis, they prove the existence of pure strategy Nash equilibria and the computation costs, i.e., the weighted sum of delay and energy consumption, of different devices can be balanced.
@@
  • Task Scheduling: In real-world applications, tasks are not only delay-sensitive and computation-intensive but also dependent on each other, which demands sequential task processing. During such a procedure, the output of one task serves as the input for the next one, which is required to be continued until all associated tasks are completed, thereby finishing the entire task. Therefore, the above issue makes it critical to schedule the offloading of such interdependent tasks in MEC systems, so as to ensure efficient and effective application performance. To this end, the authors in [51] investigate the task offloading scheme for multiple interdependent computation tasks in MEC systems, where the interdependent tasks are modeled as a directed acyclic graph. They formulate a mixed-integer nonlinear programming problem, and propose the highest response ratio and optimization method to implement joint scheduling and offloading design, which effectively reduces the average completion time in MEC systems with interdependent tasks. Furthermore, instead of assuming all requested tasks can be completed, the authors in [94] explore the task scheduling by taking into account the impact of interdependency among different tasks on the system reliability, where a multi-priority task sequencing algorithm and a deep deterministic policy gradient-based learning algorithm are developed to minimize the system deadline violation ratio.
@@
  • Cooperative Computing: The mobile devices in MEC systems generally offload their computation-intensive tasks to edge servers for low-latency task execution while reducing communication overhead. Nevertheless, edge servers may become unavailable due to factors such as limited coverage area and exhausted computing resources, resulting in new tasks offloaded by devices not being processed instantly or failing directly. To address the above challenge, multi-user cooperative computing emerges as a promising solution to enhance the reliability of MEC systems by alleviating edge server overload and exploiting computing resources distributed among mobile devices. For instance, the authors in [52] propose a device-to-device (D2D)-assisted co-offloading framework for industrial IoT, and an efficient learning-based algorithm is developed to minimize the weighted sum of task delay and migration cost without the need of complete offloading information. To minimize the average delay under time-varying idle resources and channel conditions, the authors in [95] develop an online joint optimization for the task partitioning and parallel scheduling in a D2D-enabled MEC system, which achieves near-optimal delay performance with low computational complexity while ensuring fairness across various system states. In addition, the authors in [96] introduce a distributed multi-hop task offloading decision model for Internet of Vehicles (IoV) networks, where the peripheral vehicles with idle computing resources can be exploited to improve the task execution efficiency in dynamic vehicular environments.
@@ II-C Over-the-Air Computation
In traditional communication protocols, the signals among different devices over the same channel are deemed as co-channel interference, thus each of them demands individual and accurate reception at the receiver before the subsequent processing. In the face of emerging intelligent services, achieving functional computation of gathered data becomes the key to implementing the corresponding applications, e.g., environmental monitoring and autonomous driving, while the accuracy and timeliness of such computation determine the quality of service. In this sense, the conventional compute-after-communicate strategy may become inefficient in a large-scale network with scarce resources due to the separation of communication and computation. With this in mind, AirComp has attracted much attention to pursuing the fusion of communication and computation by harnessing the co-channel interference [97, 98, 99].
@@
Seeking to integrate communication and computation, AirComp regards the co-channel waveform superposition as a summation operation for synchronized concurrently transmitted signals, enabling compute-when-communicate manner that allows the target function to be computed during the transmission process without need for one-by-one decoding at the receiver. As all devices are required to occupy the same channel to enforce the computation for the target function, the number of radio resources utilized by AirComp is unrelated to the number of participating devices, thereby yielding a much higher spectral efficiency than conventional orthogonal multiple-access (OMA) approaches that allocate separated channels to each device. Meanwhile, compared with the non-orthogonal multiple-access (NOMA) scheme counterparts, AirComp avoids the complex design for interference cancellation and saves the time for recovering each signal in terms of the functional computation at the cost of acceptable computation errors.
@@
By implementing appropriate pre- and post-processing on transmit and receive signals, respectively, AirComp enables the receiver to compute a class of functions, namely nomographic function [100], which is defined as $f(x_{1},x_{2},\dots,x_{K})=\psi\left(\sum_{k=1}^{K}\varphi_{k}(x_{k})\right)$, where $\psi(\cdot)$ denotes the post-processing function at the receiver and $\varphi_{k}(\cdot)$ denotes the pre-processing function for data $x_{k}\in\mathbb{R}$ at device $k$, $\forall\,k\in\{1,2,\dots,K\}$. Following the expression given in the above, AirComp can be decomposed into three stages [16]: 1) transmit data pre-processing at each device; 2) summation of pre-processed data through the waveform superposition property of multiple-access channels; and 3) receive data post-processing at the access point (AP). Note that, although the signal superposition can only realize the operation of summation, AirComp is capable of computing non-linear functions [101], e.g., geometric mean and Euclidean norm, and even computing any continuous function by assembling multiple nomographic computation results [100].
@@
  • Data Fusion: Data fusion is the fundamental objective of AirComp and much of the underlying research has been conducted from this aspect [102]. For instance, the authors in [19] reveal the optimal transceiver design for AirComp under peak power constraints to minimize the distortion of data fusion in single-input single-output (SISO) networks. Meanwhile, the optimal design under sum-power constraint is obtained in [103] with closed-form solutions. By exploiting multi-antenna techniques, the authors in [16, 104, 105] develop beamforming design for AirComp and validate the performance improvement that can be achieved via diversity gain and multiplexing gain of multi-antenna arrays to reduce fusion error [104] and realize multi-modal fusion [16, 105], respectively.
@@
  • Distributed Consensus: Distributed consensus expects all devices in the network to reach a consensus on a certain message without the assistance of a central node, which can be accelerated by AirComp with its efficient data aggregation. For instance, the authors in [53] show that multiple simultaneous average operations achieved by AirComp can accelerate the convergence rate of gossip-based consensus. In addition, the authors in [106] propose beamforming design for decentralized AirComp to accelerate the distributed consensus.
@@
  • Model aggregation: Federated learning (FL) has attracted much attention in machine learning areas due to its distributed computing structure and privacy-preservation property [54]. Although transmitting model parameters reduce the communication overhead by a significant amount compared to transmitting raw data, scarce communication resources still limit the communication efficiency in wireless FL systems, which motivates the emergence of AirComp-based FL [21, 22, 107, 108]. The authors in [21] demonstrate that AirComp-based FL achieves much lower communication latency than the OFDMA-based scheme. To ensure accurate model aggregation, the authors in [22, 21, 109] develop device scheduling methods. Moreover, compression and sparsification approaches are leveraged in [110] to further reduce the communication overhead in resource-constrained wireless networks.
@@ III-B Wireless Sensing Technology
Wireless sensing, which is also called radar sensing, emits modulated radio waves, receives and processes the corresponding echo signals to complete computation tasks such as object ranging and detection, and motion recognition. Its application can be traced back to the usage of radar in the early 20th century. Because of its significant military value, radar ushered tremendous development during World War II and turned to be a key technique for empowering many major and intelligent applications, including military surveillance, remote sensing in space, and auto-driving. Based on the original continuous-wave radar, new techniques in terms of wave-form design, signal processing, and system architecture are developed to enhance its abilities regarding velocity, distance, direction, and sensitivity toward motions. In this part, the development of radar system design and the corresponding signal processing are introduced.
@@ III-B1 Radar Waveform Design and Signal Processing
  • Traditional Continuous-wave Radar: As its name suggests, traditional continuous-wave radar continuously transmits a waveform with a stable frequency and receives an echo signal from one or more targets. To detect the Doppler shift of the received echo signal, a conjugate mixing method is utilized. To be specific, both the transmit signal and the echo signal are sampled with a time interval $T$ and then transformed into the frequency domain using fast Fourier transform (FFT). By calculating the cross-correlation of the two signals and applying a low-pass filter, the frequency difference of the two signals, i.e., the Doppler frequency shift (DFS), can be extracted. Based on the DFS, the velocity can be estimated. The resolution (or the detection error) of the velocity is proportional to the sampling rate $\dfrac{1}{T}$. However, the measurement of range information i.e., distance of the target which is decided by light speed and the round-trip delay, is not available with continuous waveform. The reason is that the round-trip delay is detected by the phase shift of the echo signal, which is ambiguous to the receiver due to the lack of phase basis. That’s to say, we cannot distinguish the phase shifts $\theta$ and $(\theta+2k\pi)$ with $k$ being any integers.
  • Pulsed Wave Radar: To obtain the range information, the pulsed radar is developed, where the transmit waveform is generated by modulating a single-frequency signal by a periodic window signal with period $T$. That’s to say, the transmitter periodically produces a brief pulse of radio signal and keeps silent to receive the echo signal of the detected object. The interval between two consecutive pulses is called the inter-pulse period (IPP). The range information is detected by processing the echo signal within one transmission period, called fast-time processing. Specifically, with the time basis of the transmit signal, the exact phase shift of the two signals can be detected using their correlation in the frequency domain. Subsequently, the round-trip delay is calculated and is further used to compute the distance. It should be emphasized that IPP should be larger than the round-trip delay to make sure the echo signal of one pulse can be correctly received. The resolution of rang information is proportional to the pulse transmission period $T$. The velocity is detected using multiple pulses, called slow-time processing. With a fixed velocity, the DFS of each signal in several consecutive pulse waveforms is a single-frequency signal. Hence, by calculating the correlation of the echo signal and the transmit signal over a long time scale (i.e., over multiple transmission periods) in the frequency domain, the DFS is detected and the corresponding velocity can be calculated. The resolution of the detected DFS or velocity is proportional to $1/T$. Note that there is a tradeoff between the resolutions of range and velocity information.

### v2

@@ I-D Organization
  • Section V discusses the benefits, functions, and challenges of ISCC.
  • Section VI introduces three types of signal designs for ISCC, i.e., single-/dual-/triple-functional signal designs. Particularly, triple-functional signal design refers to using one signal to perform three functions of sensing, communication, and computation.
  • In Section VII, the network resource management paradigms for ISCC are discussed. One is joint RRM for task coexistence, which refers to resource allocation among different tasks of sensing, communication, and computation to achieve multiple goals simultaneously. The other is task-oriented ISCC resource allocation, which refers to allocating network resources among the sensing, communication, and computation modules of a task to achieve its customized goal.
  • In Section VIII, the implementation of ISCC in future advanced networks is investigated in Section VIII, including ISCC for digital-twins enabled wireless networks, computing power networks supported ISCC, SAGINs supported ISCC. And the unresolved issues in ISCC that demand future research efforts are discussed.
  • Section IX concludes the paper.
@@ II-A Overview
Emerging intelligent services over wireless networks often encompass both communication and computation processes. However, in conventional wireless networks, the communication and computation processes are separately designed, as the inherent connection between the transmitted data and subsequent service tasks is ignored. Such a separate design leads to inefficient communication and computation resource utilization and in turn degrades the execution performance of the underlying service tasks. By treating the communication and computation processes as a unified module, ICC emerges as a promising technique to enhance the system performance, where the allocation of communication and computation resources are jointly optimized to achieve a common objective, e.g., energy/delay minimization and computation accuracy maximization, thereby making full use of limited resources to improve the system performance. In this section, we provide an overview of the existing studies on ICC, focusing on the efficient design of joint communication and computation, which is illustrated by presenting two promising examples, e.g., MEC and AirComp, as shown in Fig. 3 and listed below.
@@ II-B Mobile Edge Computing
In the past few years, computation tasks over wireless networks are typically accomplished by resorting to cloud computing, which requires a large amount of data generated by devices to be collected and transmitted to a centralized server with high storage and computing capabilities, so as to perform data analysis and computation for meeting various requirements of different devices. However, with the rapid development of Internet of Things (IoT), the number of connected devices dramatically increases, leading to a severe communication burden on the wireless network and high computing pressure on the central server. As a result, cloud computing that typically relies on a remote central server is no longer suitable to support latency-critical services in future wireless networks with massive connectivity and limited radio resources. To address the above issues, MEC is emerging as a new computing paradigm thanks to the advancement of microchip computing power, which enables the edge server and mobile devices to execute computing tasks with their own computing and storage capabilities at the network edge [13].
@@
A typical MEC system consists of edge servers and mobile devices, where edge servers are generally co-located with access points (APs) and deployed in proximity to mobile devices. Herein, the mobile devices are allowed to offload the computational tasks to edge servers for executing the tasks with high computing and low latency requirements. To unleash the potential of MEC, numerous studies have explored task offloading design in MEC systems with single/multiple edge server(s) and multiple users. Specifically, the system design of MEC lies in the following aspects.
@@
  • Joint Resource Allocation: Even though the edge servers are capable of providing strong computing power to fulfill various service requirements for different devices, they still confront high computation and communication burdens when there are massive tasks from multiple devices that need to be accomplished due to the limited radio and computing resources of the MEC system. To overcome the above challenge, joint communication and computation resource allocation is required to enhance the system performance in terms of different objectives, e.g., execution delay and energy consumption. For MEC in software-defined ultra-dense wireless networks with battery-limited devices, the task placement and resource allocation can be jointly optimized to minimize the average task duration [51]. In remote areas lacking conventional edge/cloud infrastructures, a SAGIN-assisted MEC framework is developed in [52], where task scheduling and resource allocation can be jointly optimized by developing a learning-based approach to reduce the execution delay. To mitigate the communication overhead, an efficient decentralized algorithm is developed in [53] to enable task offloading in a dense wireless network with selfish mobile devices. By using game-theoretical analysis, the computation costs, i.e., the weighted sum of delay and energy consumption, of different devices can be further balanced. To efficiently support multi-user task offloading, non-orthogonal multiple access (NOMA) assisted MEC is proposed in [54], where the communication and computation resources are jointly optimized for minimizing the delay and energy consumption.
@@
  • Task Scheduling: In real-world applications (e.g., autonomous driving, augmented reality, and video surveillance), tasks are not only delay-sensitive and computation-intensive but also dependent on each other, which demands sequential task processing. Hence, the output of one task becomes the input of the next one, resulting in a chain of dependencies among different tasks. Specifically, the computation tasks can be interdependent in various forms, i.e., sequential, parallel, and a combination of both. The interdependency makes it critical to schedule the offloading of such interdependent tasks in MEC systems, so as to ensure satisfactory application performance, which can be evaluated by the overall latency and energy consumption of all tasks in the application. To this end, the interdependent tasks can be modeled as a directed acyclic graph, followed by developing a task offloading scheme for multiple interdependent computation tasks in MEC systems [55]. By formulating a mixed-integer nonlinear programming problem, a joint scheduling and offloading design is proposed to effectively reduce the average completion time of interdependent tasks in MEC systems. Furthermore, the impact of interdependency among different tasks on the system reliability can be explored to facilitate the design of task scheduling as in [56], where a multi-priority task sequencing algorithm and a deep deterministic policy gradient-based learning algorithm are developed to minimize the system deadline violation ratio. In industrial MEC networks, the assembly processing line involves multiple computing tasks with diverse priority and latency requirements. This motivates the development of hybrid NOMA and orthogonal multiple access (OMA) based task offloading [57], where the selection of NOMA and OMA is based on the communication and computation latency requirements of the underlying tasks.
@@
  • Cooperative Computing: The mobile devices in MEC systems generally offload their computation-intensive tasks to edge servers for low-latency task execution while reducing communication overhead. Nevertheless, edge servers may become unavailable due to factors such as limited coverage area and exhausted computing resources, resulting in new tasks offloaded by devices not being processed instantly or failing directly. To address the above challenge, multi-user cooperative computing emerges as a promising solution to enhance the reliability of MEC systems by alleviating edge server overload and exploiting computing resources distributed among mobile devices. For instance, a device-to-device (D2D)-assisted co-offloading framework for industrial IoT is proposed in [58], where an efficient learning-based algorithm is developed to minimize the weighted sum of task delay and migration cost without the need for complete offloading information. In D2D-enabled MEC systems with time-varying idle resources and channel conditions, the task partitioning and parallel scheduling can be jointly optimized to achieve near-optimal delay performance with low computational complexity while ensuring fairness across various system states [59]. In addition, cooperative task offloading can also be applied in Internet of Vehicles (IoV) networks, where the peripheral vehicles with idle computing resources can be exploited to improve the task execution efficiency in dynamic vehicular environments [60].
@@ II-C Over-the-Air Computation
In traditional communication systems, the signals concurrently transmitted by different devices over the same channel are deemed as co-channel interference, thus each of them demands individual and accurate reception at the receiver before the subsequent processing. In the face of emerging intelligent services, achieving functional computation of gathered data becomes the key to implementing the corresponding applications, e.g., environmental monitoring and autonomous driving, while the accuracy and timeliness of such computation determine the quality of service. In this sense, the conventional compute-after-communicate strategy may become inefficient in a large-scale network with scarce resources due to the separation of communication and computation. As a result, AirComp has attracted much attention to pursuing the fusion of communication and computation by harnessing the co-channel interference [61, 62, 63].
@@
Seeking to integrate communication and computation, AirComp regards the co-channel waveform superposition as a summation operation for synchronized concurrently transmitted signals, enabling compute-when-communicate manner that allows the target function to be computed during the transmission process without need for one-by-one decoding at the receiver. As all devices are required to occupy the same channel to enforce the computation for the target function, the number of radio resources utilized by AirComp is unrelated to the number of participating devices, thereby yielding a much higher spectral efficiency than conventional OMA approaches that allocate separated channels to each device. Although power-domain NOMA can also enable concurrent transmission over the same channel by leveraging superposition coding and successive interference cancellation, the algorithmically complex and time-consuming interference cancellation operation is inevitable before performing the subsequent function computation. As a result, compared with the NOMA counterpart, AirComp has the advantages of having low implementation complexity and saving the time for recovering each signal in terms of the functional computation at the cost of acceptable computation errors.
@@
By implementing appropriate pre- and post-processing on transmit and receive signals, respectively, AirComp enables the receiver to compute a class of functions, namely nomographic function [64], which is defined as $f(x_{1},x_{2},\dots,x_{K})=\psi\left(\sum_{k=1}^{K}\varphi_{k}(x_{k})\right)$, where $\psi(\cdot)$ denotes the post-processing function at the receiver and $\varphi_{k}(\cdot)$ denotes the pre-processing function for data $x_{k}\in\mathbb{R}$ at device $k$, $\forall\,k\in\{1,2,\dots,K\}$. Following the expression given in the above, AirComp can be decomposed into three stages [17]:
1) transmit data pre-processing at each device; 2) summation of pre-processed data through the waveform superposition property of multiple-access channels; and 3) receive data post-processing at the access point (AP). Note that, although the signal superposition can only realize the operation of summation, AirComp is capable of computing non-linear functions [65], e.g., geometric mean and Euclidean norm, and even computing any continuous function by assembling multiple nomographic computation results [64].
@@
  • Data Fusion: Data fusion is the fundamental objective of AirComp and much of the underlying research has been conducted from this aspect [66]. For instance, to minimize the distortion of data fusion in single-input single-output (SISO) networks, the optimal transceiver design for AirComp under peak power constraints is developed in [20]. Meanwhile, the optimal design under sum-power constraint is obtained in [67] with closed-form expressions. By exploiting multi-antenna techniques, the beamforming design developed in [17, 68, 69] validate the performance improvement that can be achieved via diversity gain and multiplexing gain to reduce fusion error [68] and realize multi-modal fusion [17, 69], respectively. To mitigate the reliance on perfect CSI, blind AirComp can be utilized to enable data fusion. One widely used method applies one-bit quantization at the transmitter, followed by a majority vote on the sign of the aggregated data at the receiver [70].
@@
  • Distributed Consensus: Distributed consensus expects all devices in the network to reach a consensus on a certain message without the assistance of a central node, which can be accelerated by AirComp with its efficient data aggregation. Specifically, as demonstrated in [71], multiple simultaneous average operations achieved by AirComp can accelerate the convergence rate of gossip-based consensus. Such an AirComp-assisted consensus framework is further extended to multi-cluster scenarios in [72], where the local averaging time is significantly reduced and the algorithmic convergence is established. Moreover, as demonstrated in [73], effective beamforming design is also essential to accelerate the distributed consensus of decentralized AirComp-based schemes. To facilitate accurate distributed consensus, achieving time and frequency synchronization at the receiver side is essential. To this end, timing advance [74] and AirShare [75] are effective techniques for guaranteeing time synchronization and eliminating carrier frequency offset, respectively.
@@
  • Model aggregation: Federated learning (FL) has attracted much attention in machine learning areas due to its distributed computing structure and privacy-preservation property [76]. Although transmitting model parameters reduce the communication overhead by a significant amount compared to transmitting raw data, scarce communication resources still limit the communication efficiency in wireless FL systems, which motivates the emergence of AirComp-based FL [22, 23, 77, 78]. As demonstrated in [22], AirComp-based FL achieves much lower communication latency than the OFDMA-based scheme. Device scheduling is an important issue for ensuring accurate model aggregation for FL [23, 22, 79]. Moreover, compression and sparsification can also be integrated with AirComp to reduce the communication overhead of FL in resource-constrained wireless networks, as illustrated in [80]. AirComp can also be utilized to support multiple FL tasks in multi-cell wireless networks, where co-channel interference degrades the accuracy of model aggregation. This issue can be alleviated by developing a cooperative optimization framework that balances the learning performance across different FL tasks [77].
II-D Discussion
This section presents the principles and potentials of ICC, with a specific emphasis on elaborating two representative ICC techniques, i.e., MEC and AirComp. For both MEC and AirComp, the joint optimization of communication and computation resources is critical for enhancing the system performance. Moreover, task scheduling and cooperative computing are two key aspects of MEC, while the transceiver design for different scenarios is essential for AirComp. Various optimization-based algorithms have been developed to achieve efficient resource utilization in ICC. As 6G networks are expected to feature more complex architectures and larger network sizes, learning-based algorithms that have the potential to reduce computation complexity and enhance scalability for the joint communication and computation resource allocation deserve further investigation.
@@ III-B Wireless Sensing Technology
Wireless sensing, which is also called radar sensing, emits modulated radio waves, receives and processes the corresponding echo signals to complete computation tasks such as object ranging and detection, and motion recognition. Based on the original continuous-wave radar, new techniques in terms of wave-form design, signal processing, and system architecture are developed to enhance its abilities regarding velocity, distance, direction, and sensitivity toward motions. In this part, the development of radar system design and the corresponding signal processing are introduced.
@@ III-B1 Radar Waveform Design and Signal Processing
  • Traditional Continuous-wave Radar: This type of radar continuously transmits a waveform with a stable frequency and receives an echo signal from one or more targets. To be specific, both the transmit signal and the echo signal are sampled with a time interval $T$ and then transformed into the frequency domain using fast Fourier transform (FFT). By calculating the cross-correlation of the two signals and applying a low-pass filter, the Doppler frequency shift (DFS) can be extracted for estimating the velocity. The resolution (or the detection error) of the velocity is proportional to the sampling rate $1/T$. However, the measurement of range information i.e., distance of the target which is decided by light speed and the round-trip delay, is not available with continuous waveform, because the phase shift of the echo signal is ambiguous to the receiver due to the lack of phase basis. That’s to say, we cannot distinguish the phase shifts $\theta$ and $(\theta+2k\pi)$ with $k$ being any integer.
  • Pulsed Wave Radar: To obtain the range information, the pulsed radar is developed, where the transmit waveform is generated by modulating a single-frequency signal by a periodic window signal with period $T$. That’s to say, the transmitter periodically produces a brief pulse of radio signal and keeps silent to receive the echo signal of the detected object. The interval between two consecutive pulses is called the inter-pulse period (IPP). The range information is detected by processing the echo signal within one transmission period, called fast-time processing. The resolution of rang information is proportional to the pulse transmission period $T$. The velocity is detected using multiple pulses, called slow-time processing. With a fixed velocity, the DFS of each signal in several consecutive pulse waveforms is a single-frequency signal. Hence, by calculating the correlation of the echo signal and the transmit signal over a long time scale (i.e., over multiple transmission periods) in the frequency domain, the DFS is detected and the corresponding velocity can be calculated. The resolution of the detected DFS or velocity is proportional to $1/T$. Note that there is a tradeoff between the resolutions of range and velocity information.